• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sarty

    Log entry time 00:16:27 on July 26, 2006

    Entry number 175263

    keyword=Shift summary (production 90 deg: Kin 1, 2, 3)

    16:30 Beam back! Start with several checks of raster size (see several halogs). Beam spot size was checked by a harp scan done by MCC who said on the phone it was about 0.3 x 0.4 mm. Have to locate an ELOG that says that (but couldn't find one). Also steering is not great, but liveable (see halog entry 175199 for direct BPM numbers).

    17:55 Lead target in, ready to start production! 20 uA on target. Raster set to the new nominal "10 x7.5" (that's the request to MCC). Emily is doing now a quick rough setting of BCM calibration constants on the Hall A Current Monitor Screen (since just at this moment Hall C has no beam - they are in the midst of a BCM calibration sequence), making sure the coefficiencts (which will get recorded in start-run file) give the same readback for the up/down stream BCMs as OLO2. They are roughly set in, will do full calibration later.

    18:15 After 1 production run: have to stop for an access because Seonho and Alex indicate there is a TDC problem that needs attention. This is fixed by replacing a defective Lecroy 4418 delay module.

    19:45 Back to Beam Permit. At JP's suggestion, we reduce raster size to make something more like 3 mm x 3 mm. Request now "6 x 4.5".

    20:25 LD2 target in (already finished Pb point). Continue along with smooth production runs: (see Run summary below) Finish Kin 1 @ 21:56 21:56 - 22:04 Lower HRS-L Momentum for Kin 2 Start Kin 2. Finish Kin 2 @ 23:10 23:10 - 23:19: Lower HRS-L Monetum for Kin 3 23:22: Start Kin 3

    ******************************************* Run Summary:

    Run 1648: Pb target; prescale=3 gives ~300 Hz DAQ rate, 0% deadtime JUNK (TDC problem)

    After back up: to maximize DAQ rate with only ~5% deadtime current is reduced during run 1651 from 20 down to 11.5 uA

    Kin 1 Production (90 deg):

    Run 1652: Pb target: prescale=1, I=12 uA, 500 Hz DAQ< ~5-6% deadtime 600 k events (1st perfect production run!!) Run 1653: LD2 target: prescale=1, I=20 uA, 200 Hz DAQ, ~2% deadtime 152 k events Run 1654 (L) and 2230 (R): Ta target (16k and 21k evetns) 20uA Run 1655 (L) and 2231 (R): C slanted target (539k events quick in 2231) 87 k events Run 1656 (L) and 2232 (R): Fe slanted target (1M events quick in 2232)

    Kin 2 Production (90 deg):

    Run 1657: Fe slanted target (166 k events) Run 1658: C slanted (156 k) Run 1659: LD2 ... short run with too high deadtime (~40%) JUNK Run 1660: LD2 - reduced I=9 uA...600 Hz DAQ rate, only 1% deadtime(?) 175 k events Run 1661: Pb - reduced I=8uA...1000 Hz DAQ rate, 4% deadtime, presc=3 648 k events

    Kin 3 Production (90 deg):

    Run 1662: Pb...current back to I=18 uA, presc=2, ~900 Hz DAQ (6% dead) 650 k Run 1663: LD2 Run 1664: C

    This run is going as shift ends.